# Simulation-based inference using surjective sequential neural likelihood estimation

### v2

@@ 4 Experiments
We compare SSNL to SNL, SNPE-C and SNRE-C (Papamakarios et al., 2019; Greenberg et al., 2019; Miller et al., 2022) on seven synthetic experiments, a solar dynamo model from the astrophysics literature and a neural mass model from neuroscience to highlight the advantages and disadvantages of the method. We chose SNPE-C as neural posterior method since we found it is still state-of-the-art or at least highly competitive on a large number of experimental benchmarks (see, e.g., Deistler et al. (2022); Wildberger et al. (2023)). Similarly, SNRE-C is to our knowledge state-of-the-art among methods for neural ratio estimation.
We followed the experimental details of Papamakarios et al. (2019), Greenberg et al. (2019) and Miller et al. (2022). In short, for SSNL, SNL and SNPE-C, we use MAF normalizing flows as density estimators consisting of five layers (Papamakarios et al., 2017) with permutation layers in between. If not noted otherwise, each layer uses neural network with two hidden layers and $50$ nodes per layer (Germain et al., 2015). SSNL uses a dimensionality-reducing surjection in the middle layer for which the conditional density $p\left(z_{k}^{-}|f^{-1}_{k}(z_{k}^{+};z_{k}^{-},\theta),\theta\right)$ is parameterized using an MLP with two layers of $50$ nodes each. We selected the third layer as surjection such that the entire data set is "processed" once in each direction before reducing the dimensionality of the data. We evaluated surjection layers that reduce the dimensionality by $25\%$, $50\%$ or $75\%$, respectively. For SNRE-C, we use a classifier MLP that uses two layers of $50$ nodes each (see Appendix D for exact experimental details).
For each experimental model, we sample a vector of true parameters $\theta_{0}\sim p(\theta)$ and then simulate an observation $y_{0}\leftarrow sim(\theta_{0})$ which is then used to approximate $p(\theta|y_{0})$. We repeat this data generating process for $10$ different seeds. We evaluate each method sequentially in $R=15$ rounds: in each round $r$ we draw a sample $\{\theta^{r}_{n}\}_{n=1}^{N}$ of size $N=1000$ from the trained surrogate posterior (or prior if in the first round, respectively), simulate observations $y^{r}_{n}\leftarrow\text{sim}(\theta^{r}_{n})$, and train the density estimator/classifier on all available data (i.e., including from all previous rounds). After training, we compare samples from the posterior approximation of a method of each round to samples obtained from MCMC and compute divergence measures between the two samples. For the solar dynamo and neural mass model, we compare the surrogate posterior samples to the true parameter values $\theta_{0}$ as in prior work, e.g., Rodrigues et al. (2021) or Buckwar et al. (2020).
@@ 4.1 Comparing posterior distributions
Figure 1: Method accuracies on SLCP with four different evaluation measures (lower is better; y-axis shows measure).
Previous work has evaluated the accuracy of the approximated posteriors to the true posterior (or rather the posterior obtained via Monte Carlo samples), mainly using maximum mean discrepancy (MMD; Gretton et al. (2012); Sutherland et al. (2017)) and classifier two-sample tests (C2ST; Lopez-Paz & Oquab (2017)). Recently, Zhao et al. (2022) introduced a general H-divergence to assess the similarity of two (empirical) distributions, $p$ and $q$, and demonstrated that their method has higher power than members of the MMD and C2ST families in several experimental evaluations while having a low number of hyperparameters to optimize. Specifically, Zhao et al. (2022) propose to use the divergence
@@
where the H-min divergence $D^{\text{Min}}_{\ell}=H_{\ell}\left(\tfrac{p+q}{2}\right)-\text{min}(H_{\ell}(p),H_{\ell}(q))$ and H-Jensen Shannon divergence $D^{\text{JS}}_{\ell}=H_{\ell}\left(\tfrac{p+q}{2}\right)-\frac{1}{2}(H_{\ell}(p),H_{\ell}(q))$ are special cases. $H_{\ell}(p)=\text{inf}_{a\in\mathcal{A}}\mathbb{E}_{p}[\ell(X,a)]$ is the Bayes optimal loss of some decision function over an action set $\mathcal{A}$ and the loss $\ell$ can in practice be implemented, e.g., using the negative log-likelihood of a density estimator such as kernel density estimator or Gaussian mixture model (see Appendix D for details on H-divergence estimation).
We evaluated H-Min and H-Jensen Shannon divergences on the notorious simple-likelihood-complex-posterior model (SLCP; see Appendix E for a description) following the experimental details in Zhao et al. (2022) and compared them to MMD and C2ST distances (Figure 1). We found that the profiles of H-Min or H-Jensen Shannon have similar trends as C2ST and MMD, respectively (the H-Jensen Shannon divergence is in fact strictly larger than the family of MMD distances (Zhao et al., 2022)).
Hence, we propose to use both the H-Min and H-Jensen Shannon divergences as model evaluation metrics due to their power, implementational simplicity and low number of tunable hyperparameters, and will report them for the experimental evaluations. Note that in the SLCP example, which we used to assess the different divergences, SSNL outperforms the three baselines consistently with sufficiently large sample size.
@@ 4.2 SBI model benchmarks
We first evaluate SSNL on several common benchmark models from the SBI literature (Ornstein-Uhlenbeck, Lotka-Volterra, SIR, GLM) and show that it achieves state-of-the-art performance. We then demonstrate on two negative examples where SSNL, expectedly, fails to outperform the baselines (Gaussian mixture model and hyperboloid model). We first describe the experimental models briefly and then illustrate the results (see Appendix E for exact details on experimental models).
Ornstein-Uhlenbeck
The Ornstein-Uhlenbeck (OU) process (Särkkä & Solin, 2019) is a stochastic differential equation (SDE) of the form
$dY_{t}=-\theta_{2}(Y_{t}-\theta_{1})\mathrm{d}t+\theta_{3}\mathrm{d}W_{t}$
where $W_{t}$ is a Wiener process and $\theta$ are the parameters of interest. Each $Y_{t}$ above can be sampled from analytically using
We set $\theta_{1}\sim\mathcal{U}(0,10)$, $\theta_{2}\sim\mathcal{U}(0,5)$ and $\theta_{3}\sim\mathcal{U}(0,2)$, and initialize $y_{0}=0$. We then sample $100$ equally-spaced observations $Y_{t}$ where $t\in[0,10]$.
Lotka-Volterra
The Lotka-Volterra (LV) model is an ecological model that describes the dynamics of predator-prey populations:
$\begin{split}\tfrac{\mathrm{d}X_{1}}{\mathrm{d}t}&=\theta_{1}X_{1}-\theta_{2}X_{1}X_{2}\\ \tfrac{\mathrm{d}X_{2}}{\mathrm{d}t}&=-\theta_{3}X_{1}+\theta_{4}X_{1}X_{2}\\ (Y_{t1},Y_{t2})&\sim\text{LogNormal}\left(\log\left(X_{t1},X_{t2}\right),0.1\right)\\ \end{split}$
We follow the parameterization in Lueckmann et al. (2021) and set $\theta_{1}\sim\text{LogNormal}(-0.125,0.5)$, $\theta_{2}\sim\text{LogNormal}(-3,0.5)$, $\theta_{3}\sim\text{LogNormal}(-0.125,0.5)$, $\theta_{4}\sim\text{LogNormal}(-3,0.5)$, and $y_{0}=[30,1]^{T}$. We sample 50 equally-spaced observations $Y_{t}$ where $t\in[0,30]$.
SIR model
The SIR model is an epidemiological model that describes the dynamics of the number of individuals in three compartmental states: susceptible, infectious, or recovered. We adopt the presentation by Lueckmann et al. (2021):
$\begin{split}\tfrac{\mathrm{d}S}{\mathrm{d}t}&=-\theta_{1}\tfrac{SI}{K},\quad\tfrac{\mathrm{d}I}{\mathrm{d}t}=\theta_{1}\tfrac{SI}{K}-\theta_{2}I,\quad\tfrac{\mathrm{d}R}{\mathrm{d}t}=\theta_{2}I\\ Y_{t}&\sim\text{Binomial}\left(1000,\tfrac{I_{t}}{K}\right)\end{split}$
[…]
Generalized linear model
We also evaluated the methods on a Beta generalized linear regression model (Beta GLM). [^1] In Lueckmann et al. (2017), it is instead proposed to evaluate a $100$-dimensional Bernoulli GLM (which they label ”Bernoulli GLM raw”). Since SBI methods that approximate the likelihood and not the posterior, such as SSNL and SNL, can generally not model discrete data sets adequately, we instead generate continuous bounded data. Our generative model follows the exposition by Lueckmann et al. (2021), but replaces the Bernoulli likelihood with a Beta likelihood. Specifically:
$\begin{split}\eta&=X\theta,\quad\mu=\text{sigmoid}(\eta)\\ Y&\sim\text{Beta}(\mu c,(1-\mu)c)\end{split}$
where $\theta\sim\mathcal{N}(0,B)$, $X$ is a fixed design vector, and $B$ is a precomputed covariance matrix (see Appendix E for details) and $c\in\mathbb{R}^{+}$ is a non-negative concentration parameter.
@@ Results
For SSNL, we first determined the optimal embedding dimensionality in the following way: we extract the validation loss, i.e., the negative log-likelihood on the validation set, after training and use the embedding dimensionality corresponding to the network that achieved the lowest validation loss (Figure 2b) [^2] This could be done more rigorously, e.g., by splitting the data into an additional test set and evaluating its loss or by simply reducing the dimensionality sequentially such that the embedding has as many dimensions as required summary statistics, but we found this simple heuristic to be sufficient and it does not require additional computation. Furthermore, since the data is simulated with iid noise, the likelihoods on validation and test sets should be almost equal.. Since the loss profiles on all four experiments are roughly the same for each parameterization, we, for simplicity, chose to use the networks that reduce the dimensionality to $75\%$ for each experimental model.
SSNL consistently outperforms SNL, SNPE-C and SNRE-C on Ornstein-Uhlenbeck and Lotka-Volterra w.r.t. H-Min divergence, and is on par with SNL on SIR and Beta GLM (see Figure 2a). The SIR model is the only case with mixed, inconsistent results where for different rounds SNL gets outperformed by SSNL or vice versa. We assessed in which case and why SSNL has a performance advantage over SNL and argue that the autocorrelation and intrinsic dimensionality of a data set might be indicative of it (see Figure 2c). For the Ornstein-Uhlenbeck process the autocorrelation converges to zero when considering longer lags meaning that information beyond a certain point is not informative of the parameters any more favouring dimensionality reducing methods. Similarly, for the Lotka-Volterra process the autocorrelation patterns are repetitive after a certain lag meaning that the data at larger time points is basically a copy of previous time points (compare appendical Figure 7). In the case of the SIR model, the autocorrelation has first an downward trend and then changes the sign of its gradient function after reaching a saddle-point. In this case the entire time-series is informative of the parameters and dimensionality reduction has supposedly only little advantage over dimension-preservation (more experimental results can be found in Appendix F).
@@
Figure 2: Experimental results of OU, LV, SIR and GLM models. (a) H-Min divergences of the four models plotted against the accumulated data size (lower is better). SSNL is outperforming the baselines consistently in three out of four tasks, and has a particularly significant advantage in the OU and LV models. (b) Validation likelihood profiles of SSNL models when the middle layer reduces the data dimensionality by $25\%$, $50\%$ or $75\%$, respectively. The performances are similar for all models which is why we used the most conservative reduction, i.e. $75\%$ for all models (same x-axis as in (a)). (c) Autocorrelation (AR) plots for the three time series models up to a lag of $40$ (on the x-axis). The AR for the OU model converges to zero while the AR for the LV has a self-repeating structure. The SIR model has a single saddle-point and does not converge.
@@
The benefit of dimension reduction depends arguably on the length of the time series and its signal-to-noise ratio. To validate this hypothesis, we replicated the Ornstein-Uhlenbeck and Lotka-Volterra experiments but increased the dimensionality from $100$ to $1000$ and observed that the performance difference between SSNL and SNL in fact increases. For OR, SSNL has a significant performance advantage over SNL, while for LV the same can be observed with sufficient sample size (Figure 3). The requirement for a slightly higher sample size is possibly due to learning the high-dimensional conditional density $p\left(y_{-}|f^{-1}(y_{+};y_{-},\theta)\right)$.
@@
The performance of SSNL depends on whether the parameter-related information in the data can be represented in a lower-dimensional space. In scenarios where this is not the case, e.g., on Gaussian mixture or hyperboloid models, SNPE-C or SNL expectedly outperform SSNL (see appendical Figure 10).
@@ 4.3 Solar dynamo
We applied SSNL to a real-world solar dynamo model from the solar physics literature that models the magnetic field strength of the sun (see Charbonneau et al. (2005) and references therein). The model is a non-linear time series model with both additive and multiplicative noise terms
@@
$\alpha_{i}$ $\sim\mathcal{U}(\theta_{1},\theta_{1}+\theta_{2}),\quad\epsilon_{i}\sim\mathcal{U}(0,\theta_{3}),$
$y_{t+1}$ $\leftarrow\alpha_{i}g(y_{t})y_{t}+\epsilon_{i}$
The model is interesting, because it has more noise components than observed outcomes and integrating out the noise components yields a marginal likelihood that is outside the exponential family. Consequently, the number of sufficient statistics for such a model is unbounded (with the length of the time series $T$) according to Pitman-Koopman-Darmois theorem. We choose the (hyper)parameters as in Albert et al. (2022), i.e., $\theta_{1}\sim\mathcal{U}(0.9,1.4)$, $\theta_{2}\sim\mathcal{U}(0.05,0.25)$ and $\theta_{3}\sim\mathcal{U}(0.02,0.15)$, and simulate a single time series of length $T=100$ (see Appendix E.8 for details).
SSNL consistently outperforms the three baselines for this experiment (Figure 5 left column). Having a closer look at the posterior distributions of one experimental run, one can observe that SSNL already after the first round recovers the true parameters reliably while the posterior mean of SNL is heavily biased (Figure 4). After the final round, both methods converge to the true parameter values.

### v3

@@ 4 Experiments
Figure 2: Method accuracies on SLCP with different evaluation measures (lower is better; y-axis shows measure).
We compare SSNL to Sequential Neural Likelihood (SNL, Papamakarios et al., (2019)), Sequential Neural Posterior Estimation-C (SNPE-C, Greenberg et al., (2019)), Sequential Neural Ratio Estimation-C (SNRE-C, Miller et al., (2022)), Sequential Neural Approximate Sufficient Statistics (SNASS, Chen et al., (2021) and Sequential Neural Approximate Slice Sufficient Statistics (SNASS, Chen et al., (2023)) on seven synthetic experiments, a solar dynamo model from the astrophysics literature and a neural mass model from neuroscience to highlight the advantages and disadvantages of the method. We chose SNPE-C as neural posterior method since we found it is still state-of-the-art or at least highly competitive on a large number of experimental benchmarks (see, e.g., Deistler et al., (2022); Wildberger et al., (2023)). Similarly, SNRE-C is to our knowledge state-of-the-art among methods for neural ratio estimation. Conceptually related to our method, SNASS and SNASSS first compute a set of near-sufficient summary statistics using embedding networks and then use SNL to fit a posterior approximation.
We followed the experimental details of Papamakarios et al., (2019), Greenberg et al., (2019) and Miller et al., (2022). In short, for SSNL, SNL and SNPE-C we use masked autoregressive flows (MAFs) with five flow layers where each layer uses a neural network with two hidden layers and $50$ nodes per layer. Since SNASS and SNASSS have to fit additional summary and critic networks, these use MAFs with three and two layers, respectively, to have roughly the same number of parameters as the previous methods. SSNL uses a dimensionality-reducing surjection in the middle layer for which the conditional density $p\left(z_{k}^{-}|f^{-1}_{k}(z_{k}^{+};z_{k}^{-},\theta),\theta\right)$ is parameterized using an MLP with two layers of $50$ nodes each. We selected the third layer as surjection such that the entire data set is "processed" once in each direction before reducing the dimensionality of the data. We evaluated surjection layers that reduce the dimensionality by $25\%$, $50\%$ or $75\%$, respectively (see Appendix D for all experimental details).
[…]
@@ 4.1 Comparing posterior distributions
Previous work has evaluated the accuracy of the approximated posteriors to the true posterior (or rather the posterior obtained via Monte Carlo samples), mainly using maximum mean discrepancy (MMD; Gretton et al., (2012); Sutherland et al., (2017)) and classifier two-sample tests (C2ST; Lopez-Paz and Oquab, (2017)). Recently, Zhao et al., (2022) introduced a general H-divergence to assess the similarity of two (empirical) distributions, $p$ and $q$, and demonstrated that their method has higher power than members of the MMD and C2ST families in several experimental evaluations while having a low number of hyperparameters to optimize. Specifically, Zhao et al., (2022) propose to use the divergence
@@
where the H-min divergence $D^{\text{Min}}_{\ell}=H_{\ell}\left(\tfrac{p+q}{2}\right)-\text{min}(H_{\ell}(p),H_{\ell}(q))$ and H-Jensen Shannon divergence $D^{\text{JS}}_{\ell}=H_{\ell}\left(\tfrac{p+q}{2}\right)-\frac{1}{2}(H_{\ell}(p),H_{\ell}(q))$ are special cases. $H_{\ell}(p)=\text{inf}_{a\in\mathcal{A}}\mathbb{E}_{p}[\ell(X,a)]$ is the Bayes optimal loss of some decision function over an action set $\mathcal{A}$ and the loss $\ell$ can in practice be implemented, e.g., using the negative log-likelihood of a density estimator such as kernel density estimator or Gaussian mixture model (see Appendix D for details on H-divergences).
We evaluated H-Min and H-Jensen Shannon divergences on the notorious simple-likelihood-complex-posterior model (SLCP; see Appendix E for a description) following the experimental details in Zhao et al., (2022) and compared them to MMD and C2ST distances (Figure 2). We found that the profiles of H-Min or H-Jensen Shannon have similar trends as C2ST and MMD, respectively (the H-Jensen Shannon divergence is in fact strictly larger than the family of MMD distances (Zhao et al.,, 2022)).
Hence, we propose to use both the H-Min and H-Jensen Shannon divergences as model evaluation metrics for SBI benchmarks due to their power, implementational simplicity and low number of tunable hyperparameters, and will report them for the experimental evaluations. Note that in the SLCP example, which we used to assess the different divergences, SSNL outperforms the three baselines consistently with a sufficient simulation budget.
@@ 4.2 SBI model benchmarks
We first evaluate SSNL on multiple benchmark models from the SBI literature (i.e., Ornstein-Uhlenbeck, Lotka-Volterra, SIR, and generalized linear model (GLM)) and discuss when it should have performance benefits over alternative methods. We then demonstrate using two negative examples where SSNL breaks and where it should fail to outperform the baselines (Gaussian mixture model and hyperboloid model). For a detailed description of the six experimental models which we omit here, we refer to Appendix E.
@@ Results
For SSNL, we first determined the optimal embedding dimensionality in the following way: we extract the validation loss, i.e., the negative log-likelihood on the validation set, after training and use the embedding dimensionality corresponding to the network that achieved the lowest validation loss (Figure 3b) [^1] This could be done more rigorously, e.g., by splitting the data into an additional test set and evaluating its loss or by simply reducing the dimensionality sequentially such that the embedding has as many dimensions as required summary statistics, but we found this simple heuristic to be sufficient and it does not require additional computation. Furthermore, since the data is simulated with iid noise, the likelihoods on validation and test sets should be almost equal. Alternatively, information-theoretical approaches could also be applied.. Since the loss profiles on all four experiments are roughly the same for each parameterization, we, for simplicity, chose to use the networks that reduce the dimensionality to $75\%$ for each experimental model.
For the two time series models Ornstein-Uhlenbeck (OU) and Lotka-Volterra (LV), SSNL consistently outperforms all baselines. SSNL is on par with SNL on the SIR and Beta GLM models (see Figure 3a). The SIR model is the only case with mixed, inconsistent results where for different simulation budgets SNL gets outperformed by SSNL or vice versa (the figures do not show SNPE-C for LV and SIR due to its bad performance, see Figure 9 in the Appendix for complete results).
@@
Figure 3: Experimental results of OU, LV, SIR and GLM models. (a) H-Min divergences of all models plotted against the size of simulated data (lower is better). (b) Validation likelihood profiles of SSNL models when the middle layer reduces the dimensionality by $25\%$, $50\%$ or $75\%$, respectively. The performances are similar for all models which is why we used the most conservative reduction, i.e. $75\%$ for all models. (c) Autocorrelation (AR) plots for the three time series models up to a lag of $40$ (black shades correspond to realizations of a time series with different parameter values). The AR for the OU model converges to zero while the AR for the LV has a self-repeating structure. The SIR model has a single saddle-point and does not converge.
Autocorrelation and intrinsic dimensionality
We assessed in which case and why SSNL has a performance advantage over SNL and argue that a combination of autocorrelation and intrinsic dimensionality (ID) of a data set might be indicative of it (see Figure 3c where realizations of a time series model with different parameter values are shown). Notably, for the Ornstein-Uhlenbeck and Lotka-Volterra models the autocorrelation pattern seems to benefit dimensionality-reducing methods. For the Ornstein-Uhlenbeck process, the autocorrelation converges to zero when considering longer lags meaning that information beyond a certain point is not informative of the parameters any more. Similarly, for the Lotka-Volterra process the autocorrelation patterns are repetitive after a certain lag meaning that the data at larger time points is basically a copy of previous time points (compare Figure 7 in the Appendix). In the case of the SIR model, the autocorrelation has first an negative slope and then changes the sign of its gradient function after reaching a saddle-point. Consequently, the entire time-series is informative of the parameters and dimensionality reduction has supposedly only little advantage over dimension-preservation (more experimental results can be found in Appendix F).
@@
The benefit of dimension reduction depends arguably on the length of the time series and its signal-to-noise ratio. To validate this hypothesis, we replicated the OU and LV experiments but increased the number of time points from $100$ to $1000$ and observed that the performance difference between SSNL and SNL in fact increases. For Ornstein-Uhlenbeck, SSNL has a significant performance advantage over SNL, while for Lotka-Volterra the same can be observed with sufficient simulation budget (Figure 4). We hypothesize that this is due to the fact that in some cases learning the high-dimensional conditional density $p\left(y_{-}|f^{-1}(y_{+};y_{-},\theta)\right)$ requires an increased sample size.
@@
The performance of SSNL depends on whether the parameter-related information in the data can be represented in a lower-dimensional space. In scenarios where this is not the case, e.g., on Gaussian mixture or hyperboloid models, SNPE-C or SNL expectedly outperform SSNL (see Figure F.2 in the Appendix).
@@ 4.3 Solar dynamo
We applied SSNL to a real-world solar dynamo model from the solar physics literature that models the magnetic field strength of the sun (see Charbonneau et al., (2005) and references therein). The model is a non-linear time series model with both additive and multiplicative noise terms
@@
$\alpha_{t}$ $\sim\mathcal{U}(\theta_{1},\theta_{1}+\theta_{2}),\quad\epsilon_{t}\sim\mathcal{U}(0,\theta_{3}),$
$y_{t+1}$ $\leftarrow\alpha_{t}g(y_{t})y_{t}+\epsilon_{t}$
The model is interesting, because it has more noise components than observed outcomes and integrating out the noise components yields a marginal likelihood that is outside the exponential family. Consequently, the number of sufficient statistics for such a model is unbounded (with the length of the time series $T$) according to Pitman-Koopman-Darmois theorem. We choose the prior $p(\theta)$ and hyperparameters $b$ and $w$ as in Albert et al., (2022) and simulate a single time series of length $T=100$ (see Appendix E.8 for details).
SSNL consistently outperforms the five baselines for this experiment (Figure 5a left column). Having a closer look at the posterior distributions of one experimental run, one can observe that SSNL already after the first round recovers the true parameters reliably while the posterior mean of SNL is heavily biased (Figure 6). After the final round, both methods converge to the true parameter values.
4.4 Neural mass model
We also evaluate SSNL on the stochastic version of the Jansen-Rit neural mass model (Ableidinger et al.,, 2017) which describes the collective electrical activity of neurons by modelling interactions of cells (see Ableidinger et al., (2017); Buckwar et al., (2020); Rodrigues et al., (2021) for details). The model is a $6$-dimensional SDE of the form
$\begin{split}\mathrm{d}\begin{pmatrix}R_{t}\\ S_{t}\end{pmatrix}=\begin{pmatrix}S_{t}\\ -\Gamma^{2}R_{t}-2\Gamma S_{t}+G_{\theta}(R_{t})\end{pmatrix}\mathrm{d}t+\begin{pmatrix}0\\ \Sigma_{\theta}\end{pmatrix}\mathrm{d}W_{t}\end{split}$
where $R=[Y_{1},Y_{2},Y_{3}]^{T}$, $S=[Y_{4},Y_{5},Y_{6}]^{T}$, $W_{t}$ is a Wiener process, $\Sigma_{\theta}$ is a diagonal covariance matrix, $G_{\theta}$ is a displacement vector, $\Gamma$ is a matrix, and $\theta$ is a four-dimensional random vector with uniform prior (see Appendix E.9 for details).
With a sufficient simulation budget, in this case $4000$ simulations, SSNL convincingly outperforms the baselines having the lowest MSE. As before (see, e.g., Figure 4), we hypothesize that the performance of SSNL is worse for lower simulation budgets, since an additional conditional density has to be learned. Intriguingly, the inferences of SNL, which like SSNL approximates the likelihood function, are significantly worse than for SSNL and overall inconsistent, indicating that SSNL is in general an excellent off-the-shelf estimator for high-dimensional data sets (more experimental results can be found in Appendix F).
(a) Solar dynamo
(b) Neural mass model
Figure 5: Solar dynamo and neural mass model evaluation (we show the MSE w.r.t. the prior sample $\theta_{\text{obs}}$ that was used to simulate the observation $y_{\text{obs}}$. Values are normalized by the minimum MSE in the entire data set).